OpenStudy (anonymous):

Show that A={(x,y)R^2/ 1

OpenStudy (anonymous):

Take a=(x,y) in A. Let \[ c =\sqrt{x^2+y^2} \\ r=\frac 1 2 Minimum\{ c-1, 7-c\} \] The disk of center a and radius r is contained in A. We showed that every element in a belongs to disk that is contained in A. So A is open.